A plumber makes 60 an hour if he already has saved 65 how Many hours must he work to save a least 275
Write,solve and check inequality
1
Expert's answer
2013-03-13T08:42:54-0400
Suppose, he must work X hours to save a least 275:
60*X + 65 >= 275
60*X >= 275 - 65 = 210
X >= 210/60
X>= 3.5
For X = 3.5 he save:
60*3.5 + 65 = 210 + 65 = 275
Therefore
Answer: he must work (to save a least 275) 3.5 orlonger hours.